OpenStudy (anonymous):

How does the ratio of the amount of radiation intercepted by the Earth compare to the amount intercepted by the Moon?

OpenStudy (anonymous):

You measure two waves with wavelengths of W_1 = 1 cm and W_2 = 0.001 cm. What is the ratio of their energies? 1.The energies have the same ratio as the wavelengths, namely, E_1 / E_2 = W_1 / W_2 = 1000. 2.The energies have the same ratio squared as the wavelengths, namely, E_1 / E_2 = ( W_1 / W_2 )^2 = 10^6. 3.The energies have the inverse squared ratio of the wavelengths, namely, E_1 / E_2 = ( W_2 / W_1 )^2 = 10^-6 4.The energies have the inverse ratio of the wavelengths, namely, E_1 / E_2 = W_2 / W_1 = 0.001.